# Studying the QCD phase diagram using pressure derivatives from lattice QCD

### 07 Mar 2026 v1

@@ 1 Introduction
Derivatives of the QCD pressure, $P$, with respect to temperature, quark masses, or the chemical potentials, $\mu_{X}$, of the conserved charges, $X$, are powerful tools for investigating the features of the QCD phase diagram [1] shown in Fig. 1. In the chiral limit of two massless quark flavors, the chiral symmetry group, $SU(2)_{L}\times SU(2)_{R}$, gets spontaneously broken down to the diagonal subgroup, $SU(2)_{V}$. At vanishing baryon chemical potential, $\mu_{B}$, this spontaneously broken chiral symmetry gets restored by a true second-order phase transition at a critical temperature $T_{c}=132^{+3}_{-6}$MeV [17]. However, the universality class of the two-flavor chiral phase transition depends upon the extent of the $U_{A}(1)$ breaking near the transition [26], and the status of the later near $T_{c}$ remains an open question. The signature of this genuine phase transition manifests at the physical value of the approximate degenerate up and down (or light) quark masses, $m_{u,d}\equiv m_{l}$, as a chiral crossover at a pseudocritical temperature, $T_{pc}=156.5\pm 1.5$ MeV [5, 11, 25]. The chiral phase transition and the chiral crossover at vanishing $\mu_{B}$ have been studied on the lattice using derivatives of the $P$ involving at least one derivative w.r.t. the symmetry breaking field, $H=m_{l}/m_{s}$, such that the strange quark mass, $m_{s}$ is always fixed to its physical value, whereas $m_{l}$ is varied from its physical value down to the chiral limit. The reason for the previous statement is that the order of the parameter of the chiral phase transition (chiral condensate) is related to the first derivative of $P$ w.r.t. $H$.
@@ 2 Universal scaling framework applied to the Taylor expansion coefficients of the pressure
Figure 2: Shown are the Taylor expansion coefficients of the QCD pressure defined in Eq.(6) for the strangeness neutral and isospin-symmetric matter. The figures are taken from Ref. [7].
@@
In the heavy-ion collisions of gold or lead, the net strangeness density, $n_{S}$, vanishes, whereas the ratio of net electric-charge density ($n_{Q}$) to net baryon number density ($n_{B}$) is close to 0.4. However, it was shown in [4], that the thermodynamic quantities do not vary much with $n_{Q}/n_{B}$, so, the isospin-symmetric case $n_{Q}/n_{B}=0.5$ is also a good approximation. Given these constraints, $\hat{\mu}_{Q}$ and $\hat{\mu}_{S}$ can be expressed as Taylor series in $\hat{\mu}_{B}$. We note that the charm quarks are formed in the initial-hard thermal collisions, so expansion in $\hat{\mu}_{C}$ is not relevant for comparison with the available experimental results. For the strangeness neutral ($n_{S}=0$) and isospin symmetric matter ($n_{Q}/n_{B}=0.5$), one obtains,
@@
For definition of $\tilde{\chi}^{B}_{2k}(T)$, see Ref. [7]. Obviously, $H$ attains its physical value in the experiments, and the results reviewed in the following will be based on $(2+1)$-flavor ensembles [8] of the HotQCD collaboration with two degenerate light and a heavier strange quark tuned to their physical mass values while keeping $m_{\ell}/m_{s}=1/27$. These data sets have been generated using the HISQ (Highly Improved Staggered Quark Action) action with tree-level coefficients and a tree-level improved gauge action. For details on the ensembles and calculation, refer to [7].
@@
In Fig. 2, the four upto $\mathcal{O}(\hat{\mu}_{B}^{8})$ Taylor expansion coeffeicients of the pressure defined in Eq. (6), follow the scaling expectation of Eq. (7). This implies that the behaviour of the energy-like observables shown in Fig. 2 can be understood using the universal scaling arguments, and at the same time, $\kappa_{2}^{B}$ can be extracted from them. For example, $P_{4}$ behaves like the specific heat that should develop a pronounced cusp in the chiral limit, but no divergence as the critical exponent $\alpha$ of a $3d$-O(N)O(N) [^3] We note that even though there are indications that the 2-flavor chiral phase transition belongs to $3d$-$O(4)$ universality class, this needs further clarification. universality class is negative with magnitude less than 1, and should receive a larger regular contribution at $T_{c}$.
@@ 3 Chiral crossover and deconfinement
$\hat{\mu}_{\ell}$ is a flavor chemical potential. For the isospin symmetric case, $\hat{\mu}_{\ell}=\hat{\mu}_{B}/3$. $\chi^{M_{\ell}}_{t(T)}$ and $\chi^{M_{\ell}}_{t(\ell,\ell)}$ are related to the derivative of the scaling function of the order parameter, $f_{G}^{\prime}(z)$, that peaks at a characteristic value $z_{t}$, whereas $\chi_{m}$ is related to the scaling function of the susceptibility of the order parameter, $f_{\chi}(z)$, that peaks at $z_{m}$ [22]. The latter is a pure magentization-like observable, while the remaining two are mixed observables. Ignoring the sub-leading corrections, different types of pseudocritical temperatures ($T_{pc,t/m}$) obtained from different types of observables are related to the critical temperature as follows:
@@
Figure 4: Shown are the relative contributins of charm quark-like excitaions $P_{q}^{C}$, charmed baryons, $P_{B}^{C}$, and charmed mesons, $P_{M}^{C}$ to the partial charm pressure, $P_{C}$. The dashed lines of the same color represent the corresponding HRG predictions. The yellow band represents $T_{pc}$ with its uncertainty. The figure is taken from Ref. [3].
@@
This agreement supports the use of hadron resonance gas (HRG) models as an effective description of strongly interacting matter near freeze-out. In the low temperature phase where HRG is valid, the fact that the Boltzmann approximation is valid for relatively heavier open-charm hadrons (as well as the charm quark), makes it easier to construct partial pressure contributions of charmed hadron pressure to QCD pressure using generalized susceptibilities involving charm (4). It was shown in Ref. [3] that the lattice results on generalized charm susceptibilities agree with their HRG counterparts in the low temperature phase, and deviate from HRG at the chiral crossover. The breakdown of HRG at $T_{pc}$ signals the onset of the melting of the charmed hadrons and appearance of charm quark-like degrees of freedom inside QGP. The non-interacting gas of charmed hadrons was extended to include charm-quark like excitations inside QGP, and was found to be valid in Ref. [3]. In this quasi-particle model, one can construct proxies for the partial pressure of charm quark, $P_{q}^{C}$, charmed baryons, $P_{B}^{C}$, and charmed mesons, $P_{M}^{C}$ as follows:
@@
As can be seen in Fig. 4, relative contributions of both $P_{B}^{C}$ and $P_{M}^{C}$ to the partial charm pressure $P_{C}\equiv\chi^{C}_{4}$ agree well with the HRG calculation, that takes into account experimentally observed charmed hadrons and the ones predicted by quark model, up to $T_{pc}$. Above $T_{pc}$, the share of charmed hadrons to $P_{C}$ starts dropping, and the relative contribution of $P_{q}^{C}$ to $P_{C}$ starts rising. $P_{q}^{C}$ becomes the dominant contributor above 175 MeV [29].These results indicate that at $\hat{\mu}_{B}=0$, deconfinement coincides with the chiral crossover. Moreover, based on the comparison of the extended HRG model explained above to the continuum lattice results [28], Ref. [21] predicted that the experimentally unobserved charmed baryons give approximately 49% contribution to the partial charmed baryon pressure at the freeze-out, whereas the charmed meson sector is experimentally well-known, and approximately 12% contribution to $P_{M}^{C}$ comes from the experimentally unknown states [27]. These results are also corroborated by the analysis of the open-charm hadron yield observed in experiments [12].
@@ 4 Taylor expansions of thermodynamic quantities and the QCD critical endpoint
Although the limit, $k\rightarrow\infty$ cannot be realized in practice, successive ratios of the available coefficients provide approximate estimates of the radius of convergence and thus yield indirect information about the analytic structure of the thermodynamic potential in the complex $\hat{\mu}_{B}$ plane.
@@
Fig. 5 [Right] maps the types of the poles in the parameter space of $c_{8,2}$ and $c_{6,2}$. In the light yellow region, denoted as cc, all four poles are complex. This same region is shown as bounded by the solid black lines in Fig. 5 [Left]. As can be seen, given the errors, except for the highest (170 MeV) and the lowest (125 MeV) temperatures, the nearest singularity to the origin is not purely real, and thus is not related to a phase transition at these temperatures. Fig. 5 [Right] also shows regions in the parameter space for the occurrence of two pairs of imaginary poles: ii; two pairs of real poles: rr; and one pair of imaginary and one pair of real poles: ir and ri, with the first letter indicating the type of pole pair closest to the origin. This means if the error on the $T=125$ MeV point could be reduced, and at the same time if the 8th order coefficient becomes positive, the green point could move out of the region of four complex poles to a region where the closest singularity to the origin is real. It is worth pointing out that the Fig. 5 [Left] is older than the bottom-right panel of Fig. 2. Notice that within errors, the eighth order coefficient is positive in Fig. 2. Ultimately, these results can be used to put bounds on the location of the QCD critical endpoint. The sketch of the QCD phase diagram in Fig. 1 depicts that $T_{c}(\hat{\mu}_{B}=0)=132^{+3}_{-6}$ MeV (the red dot) puts an upper bound on the QCD critical endpoint (the blue dot). The information extracted from a [4,4] Pade approximant is therefore consistent with this. For details on various estimates of the radius of convergence as a function of temperature, see Ref [9].

### 01 Apr 2026 v2

@@ 1 Introduction
Derivatives of the QCD pressure, $P$, with respect to temperature, quark masses, or the chemical potentials, $\mu_{X}$, of the conserved charges, $X$, are powerful tools for investigating the features of the QCD phase diagram [1] shown in Fig. 1. In the chiral limit of two massless quark flavors, the chiral symmetry group, $SU(2)_{L}\times SU(2)_{R}$, gets spontaneously broken down to the diagonal subgroup, $SU(2)_{V}$. At vanishing baryon chemical potential, $\mu_{B}$, this spontaneously broken chiral symmetry gets restored by a true second-order phase transition at a critical temperature $T_{c}=132^{+3}_{-6}$MeV [18]. However, the universality class of the two-flavor chiral phase transition depends upon the extent of the $U_{A}(1)$ breaking near the transition [27], and the status of the latter near $T_{c}$ remains an open question. The signature of this genuine phase transition manifests at the physical value of the approximate degenerate up and down (or light) quark masses, $m_{u,d}\equiv m_{l}$, as a chiral crossover at a pseudocritical temperature, $T_{pc}=156.5\pm 1.5$ MeV [5, 12, 26]. The chiral phase transition and the chiral crossover at vanishing $\mu_{B}$ have been studied on the lattice using derivatives of the $P$ involving at least one derivative w.r.t. the symmetry breaking field, $H=m_{l}/m_{s}$, such that the strange quark mass, $m_{s}$ is always fixed to its physical value, whereas $m_{l}$ is varied from its physical value down to the chiral limit. The reason for the previous statement is that the order of the parameter of the chiral phase transition (chiral condensate) is related to the first derivative of $P$ w.r.t. $H$.
@@ 2 Universal scaling framework applied to the Taylor expansion coefficients of the pressure
Figure 2: Shown are the Taylor expansion coefficients of the QCD pressure defined in Eq.(6) for the strangeness-neutral and isospin-symmetric matter. The figures are taken from Ref. [8].
@@
In the heavy-ion collisions of gold or lead, the net strangeness density, $n_{S}$, vanishes, whereas the ratio of net electric-charge density ($n_{Q}$) to net baryon number density ($n_{B}$) is close to 0.4. However, it was shown in [4] that the thermodynamic quantities do not vary much with $n_{Q}/n_{B}$, so the isospin-symmetric case $n_{Q}/n_{B}=0.5$ is also a good approximation. Given these constraints, $\hat{\mu}_{Q}$ and $\hat{\mu}_{S}$ can be expressed as Taylor series in $\hat{\mu}_{B}$. We note that the charm quarks are formed in the initial-hard thermal collisions, so expansion in $\hat{\mu}_{C}$ is not relevant for comparison with the available experimental results. For the strangeness neutral ($n_{S}=0$) and isospin symmetric matter ($n_{Q}/n_{B}=0.5$), one obtains,
@@
For the definition of $\tilde{\chi}^{B}_{2k}(T)$, see Ref. [8], and Appendix A of Ref. [6] [^3] We refer to Refs. [8, 6] for the explicit expressions of the constrained Taylor expansion coefficients, as they are rather lengthy and the discussion in this work does not rely on their detailed form.. Obviously, $H$ attains its physical value in the experiments, and the results reviewed in the following will be based on $(2+1)$-flavor ensembles [9] of the HotQCD collaboration with two degenerate light and a heavier strange quark tuned to their physical mass values while keeping $m_{\ell}/m_{s}=1/27$. These data sets have been generated using the HISQ (Highly Improved Staggered Quark Action) action with tree-level coefficients and a tree-level improved gauge action. For details on the ensembles and calculation, refer to [8].
@@
In Fig. 2, the four Taylor expansion coefficients of the pressure up to $\mathcal{O}(\hat{\mu}_{B}^{8})$, as defined in Eq. (6), follow the scaling expectation of Eq. (7). This implies that the behaviour of the energy-like observables shown in Fig. 2 can be understood using the universal scaling arguments, and at the same time, $\kappa_{2}^{B}$ can be extracted from them. For example, $P_{4}$ behaves like the specific heat that should develop a pronounced cusp in the chiral limit, but no divergence as the critical exponent $\alpha$ of a $3d$-O(N)O(N) [^4] We note that even though there are indications that the 2-flavor chiral phase transition belongs to $3d$-$O(4)$ universality class, this needs further clarification. universality class is negative with magnitude less than 1, and should receive a larger regular contribution at $T_{c}$.
@@ 3 Chiral crossover and deconfinement
$\hat{\mu}_{\ell}$ is a flavor chemical potential. For the isospin symmetric case, $\hat{\mu}_{\ell}=\hat{\mu}_{B}/3$. $\chi^{M_{\ell}}_{t(T)}$ and $\chi^{M_{\ell}}_{t(\ell,\ell)}$ are related to the derivative of the scaling function of the order parameter, $f_{G}^{\prime}(z)$, that peaks at a characteristic value $z_{t}$, whereas $\chi_{m}$ is related to the scaling function of the susceptibility of the order parameter, $f_{\chi}(z)$, that peaks at $z_{m}$ [23]. The latter is a pure magnetization-like observable, while the remaining two are mixed observables. Ignoring the sub-leading corrections, different types of pseudocritical temperatures ($T_{pc,t/m}$) obtained from different types of observables are related to the critical temperature as follows:
@@
Figure 4: Shown are the relative contributions of charm quark-like excitations $P_{q}^{C}$, charmed baryons, $P_{B}^{C}$, and charmed mesons, $P_{M}^{C}$ to the partial charm pressure, $P_{C}$. The dashed lines of the same color represent the corresponding HRG predictions. The yellow band represents $T_{pc}$ with its uncertainty. The figure is taken from Ref. [3].
@@
This agreement supports the use of hadron resonance gas (HRG) models as an effective description of strongly interacting matter near freeze-out. In the low temperature phase where HRG is valid, the fact that the Boltzmann approximation is valid for relatively heavier open-charm hadrons (as well as the charm quark), makes it easier to construct partial pressure contributions of charmed hadron pressure to QCD pressure using generalized susceptibilities involving charm (4). It was shown in Ref. [3] that the lattice results on generalized charm susceptibilities agree with their HRG counterparts in the low temperature phase, and deviate from HRG at the chiral crossover. The breakdown of HRG at $T_{pc}$ signals the onset of the melting of the charmed hadrons and appearance of charm quark-like degrees of freedom inside QGP. The non-interacting gas of charmed hadrons was extended to include charm-quark-like excitations inside QGP, and was found to be valid in Ref. [3]. In this quasi-particle model, one can construct proxies for the partial pressure of charm quark, $P_{q}^{C}$, charmed baryons, $P_{B}^{C}$, and charmed mesons, $P_{M}^{C}$ as follows:
@@
As can be seen in Fig. 4, relative contributions of both $P_{B}^{C}$ and $P_{M}^{C}$ to the partial charm pressure $P_{C}\equiv\chi^{C}_{4}$ agree well with the HRG calculation, that takes into account experimentally observed charmed hadrons and the ones predicted by quark model, up to $T_{pc}$. Above $T_{pc}$, the share of charmed hadrons to $P_{C}$ starts dropping, and the relative contribution of $P_{q}^{C}$ to $P_{C}$ starts rising. $P_{q}^{C}$ becomes the dominant contributor above 175 MeV [30].These results indicate that at $\hat{\mu}_{B}=0$, deconfinement coincides with the chiral crossover. Moreover, based on the comparison of the extended HRG model explained above to the continuum lattice results [29], Ref. [22] predicted that the experimentally unobserved charmed baryons give approximately 49% contribution to the partial charmed baryon pressure at the freeze-out. In contrast, the charmed meson sector is experimentally well-known, and approximately 12% contribution to $P_{M}^{C}$ comes from the experimentally unknown states [28]. These results are also corroborated by the analysis of the open-charm hadron yield observed in experiments [13].
@@ 4 Taylor expansions of thermodynamic quantities and the QCD critical endpoint
Although the limit, $k\rightarrow\infty$, cannot be realized in practice, successive ratios of the available coefficients provide approximate estimates of the radius of convergence and thus yield indirect information about the analytic structure of the thermodynamic potential in the complex $\hat{\mu}_{B}$ plane.
@@
Fig. 5 [Right] maps the types of the poles in the parameter space of $c_{8,2}$ and $c_{6,2}$. In the light yellow region, denoted as cc, all four poles are complex. This same region is shown as bounded by the solid black lines in Fig. 5 [Left]. As can be seen, given the errors, except for the highest (170 MeV) and the lowest (125 MeV) temperatures, the nearest singularity to the origin is not purely real, and thus is not related to a phase transition at these temperatures. Fig. 5 [Right] also shows regions in the parameter space for the occurrence of two pairs of imaginary poles: ii; two pairs of real poles: rr; and one pair of imaginary and one pair of real poles: ir and ri, with the first letter indicating the type of pole pair closest to the origin. This means that if the error on the $T=125$ MeV point could be reduced, and at the same time if the 8th order coefficient becomes positive, the green point could move out of the region of four complex poles to a region where the closest singularity to the origin is real. It is worth pointing out that the Fig. 5 [Left] is older than the bottom-right panel of Fig. 2. Notice that within errors, the eighth order coefficient is positive in Fig. 2. Ultimately, these results can be used to put bounds on the location of the QCD critical endpoint. The sketch of the QCD phase diagram in Fig. 1 depicts that $T_{c}(\hat{\mu}_{B}=0)=132^{+3}_{-6}$ MeV (the red dot) puts an upper bound on the QCD critical endpoint (the blue dot). The information extracted from a [4,4] Pade approximant is therefore consistent with this. For details on various estimates of the radius of convergence as a function of temperature, see Ref [10].
@@ 5 Conclusions
We emphasize that all results discussed in this work are taken from previously published studies. The novelty of this manuscript lies in providing a unified perspective that connects these results through derivatives of the QCD pressure, rather than in presenting new simulation data.
Acknowledgements
We acknowledge support from the DFG cluster of excellence ORIGINS, funded by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy-EXC-2094-390783311.